# AI-Powered Urban Transportation Digital Twin: Methods and Applications

### v1

@@ Abstract
We present a survey paper on methods and applications of digital twins (DT) for urban traffic management. While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived. In order to add values to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies. We will first review the DT pipeline leveraging cyberphysical systems and propose our DT architecture deployed on a real-world testbed in New York City. This survey paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
footnotetext: Xuan Di, Yongjie Fu, Mehmet K.Turkcan, and Zhaobin Mo are with the Department of Civil Engineering and Engineering Mechanics, Columbia University. Emails: {sharon.di, yf2578, mkt2126, zm2302}@columbia.edufootnotetext: Xuan Di, Zoran Kostic, and Gil Zussman are also with the Data Science Institute, Columbia University.footnotetext: Mahshid Ghasemi, Chengbo Zang, Abhishek Adhikari, Zoran Kostic, and Gil Zussman are with the Department of Electrical Engineering, Columbia University.Emails: {mg4089, cz2678, abhishek.adhikari, zk2172, gil.zussman}@columbia.edu
Keywords: Digital twin, AI, Urban transportation management, Survey
1 Introduction
@@
Urban transportation systems are complex to model and simulate, due to heterogeneous road users (such as cars, pedestrians, cyclists, scooters) interacting in multimodal traffic environments consisting of public and private travel modes. With fast-changing traffic evolution in time and space, traffic simulation if improperly calibrated, might produce traffic management strategies that largely deviate from the reality, potentially leading to suboptimal or even detrimental outcomes. With ubiquitous sensors in smart cities, it is the time to augment conventional traffic simulators, many of which were developed in an era when only “small data” became available. Emerging traffic sensors are expected to generate big volumes of data, transmitted via communication networks and processed on edge cloud computing with artificial intelligence (AI) for real-time traffic management. Such a transformation poses challenges that call for the development of a new paradigm, namely, digital twin (DT), which will push the envelope in urban transportation management.
Literally, DT is the digital replica of a physical object or asset [48], where a digital world mirrors a physical world for real-time diagnosis, prognosis, and decision making. Recent years have seen a growing amount of studies on DT in various domains, ranging from manufacturing [48], aerospace [76], to smart cities [82]. In particular, there is a sizable body of articles on transportation DTs powered by vehicular technologies such as connectivity and autonomy [134, 162, 57, 199, 80]. With such a growing trend, we would like to restrict the scope of this survey paper. Here we are particularly interested in urban traffic settings where cars have to interact with vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [153]), who are an integral part and also potential users of the DT. We will primarily focus on two use cases, namely, intersection safety warning and network wide traffic signal control, for improved traffic safety and efficiency. The literature on DTs for urban traffic management is lacking, partly because the development of a DT for a system is non-trivial, particularly involved with human participants like VRUs. The development of such a DT must engage with expertise in sensing, communication, computing, and human-centric perspectives. Accordingly, we strongly believe that a comprehensive review is warranted on trans-disciplinary topics.
@@
This paper presents a survey paper on methods and applications of DT for urban traffic management. We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods. We will propose a DT for real-time traffic monitoring and optimization [127], based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation. The overall contributions of this paper include: (1) introducing AI methods used in the DT pipeline; (2) exploring the architecture of transportation DTs and propose a prototype for reference; and (3) identifying gaps and directions.
@@
The rest of this paper is organized below. Section 2 unifies various concepts of DT and introduces the relationship between cyberphysical transportation systems and DTs; Section 3 reviews the literature on sensing and communication technologies, object detection and tracking, and urban transportation use cases. Section 4 demonstrates the architecture of our proposed DT building on a real-world testbed. Section 5 concludes our work and presents potential research directions as well as open questions.
2 Preliminaries
2.1 Variants of DT definitions
@@
The definition of DT has evolved rapidly. Despite presenting its own version, these articles all share common elements and resemble certain characteristics [48, 114]. In general, there is a physical world (aka. the physical) and a digital world (aka. the digital). The physical world evolves in time and space. To ensure that the physical system is run in a desired direction, it requires close monitoring, operation, and management. Thus, the role a digital world plays is to model and simulate the dynamics of the physical world in a synchronized fashion, so that the digital can also predict the future states of the physical precisely, which offers a ground for optimal decision making. The physical and the digital exchange data and information flows via a two-way communication. Specifically, the physical sends the data of its own state to the digital, and the digital feeds back the actuation signals to the physical. The actuation would trigger a change in the stage of the physical, and the updated state is sensed and sent back to the digital again. This iterative process runs between the physical and the digital as time unfolds. The sequential states of the physical should move towards a more desired state than without a DT. In other words, the ultimate goal of a DT development is to add values to the physical for improved safety and efficiency.
Depending on whether there is any interaction between the physical and the digital via data and/or information exchange, there are three types of digital twinning concepts [47], namely,
  1. 1.
    digital model: with no automatic data exchange;
  2. 2.
    digital shadow: with only one-way data exchange from the physical to the digital;
  3. 3.
    digital twin: with two-way communication.
Let us specify a definition of DT in transportation.
Definition 2.1
@@
Traffic digital twin is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management. It is updated online using continuously fed data collected from the physical and send control policies or issue warnings back to the physical, leveraging big data and AI tools. Traffic DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
2.2 Cyberphysical transportation systems
Definition 2.2
@@ 2.2 Cyberphysical transportation systems
2.2.1 CPS as DT technological enabler
A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig. 1). The DT encloses the cyber layer, and relies on all the modules for two-way interaction with the physical.
@@ 2.2.1 CPS as DT technological enabler
To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination. In Secs. 3.1-3.3, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec. 4.1.1.
2.2.2 DT as CPS testing tool
An automatic system like CPS demands high degree of autonomy for urban traffic management, with increased complexity and scale. Thus, DT becomes crucial for the test and verification of a CPS. We need to verify the system behavior of a CPS to ensure that it performs as desired [141], such as safety, efficiency, timeliness, and accuracy.
Due to the integration of cyber and physical components, as well as their two-way coupling via communication and networking, testing a CPS is non-trivial. Given that the digital is expected to be the twin of the physical, it is natural to compare the difference in the outputs of both the physical and the DT using “Grieves performance test” [48]. Hardware-in-the-loop testing, such as vehicle-in-the-loop [172], could help test directly the performance of some physical components “live,” while allowing the rest to be simulated within a DT. For DTs that involve the feedback to humans, human-in-the-loop test is widely used on driving simulators [183, 79]. Recent years have seen a growing trend of using augmented reality and virtual reality (AR/VR) or Metaverse to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [164, 72].
Since there does not exist a unified approach for the DT testing in the context of transportation, in Sec. 3.4, we will review evaluation metrics employed by various use cases.
3 Related work
In this section, we will review technological enablers of DT along the pipeline depicted in Fig. 1.
Figure 1: Pipeline of a CPS-structured DT.
3.1 Sensing and perception
@@
Sensors are the “eyes” (and “ears”) of a DT. Tab. 2 summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications. The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
@@
Table 1: Pros and cons of sensors for urban traffic applications (partially adapted from [154, 194])
@@
Camera Obstacle and lane detection 1. Maintains good resolution when the FOV is wide.
@@
2. Can be integrated with the vehicle system for behavior analysis [98].
@@
3.2 Object detection and tracking
Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar. Object tracking is the process of monitoring the detected objects over time to determine their position and movement. Multi-object tracking is concerned with maintaining the identity of the objects and generating their trajectories. Trajectory prediction involves forecasting the future paths of detected and tracked objects. These tasks highly rely on training datasets for urban traffic scenes. Note that there is a much larger size of public datasets collected from on-board vehicles [31], but fewer from other sensor types. Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab. 1.
Table 2: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [31].)
@@ 3.2 Object detection and tracking
Aerial VisDrone [202], NGSim [116], highD [67], rounD [68], DOTA [171], CitySim [198] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Urban spaces in China and Aacheny, intersections in California and Florida
@@
Object detection has been studied extensively for urban applications. Single-stage object detectors, following the original single shot multibox detector [93] and You Look Only Once (YOLO) [126] architectures, have become popular due to their real-time deployment capabilities. In the last few years, transformer-based object detection approaches, competitive with YOLO models, have emerged as the state-of-the-art in object detection when designed to be deployed in real-time settings [197]. Recent progress in YOLO object detection performance has been enabled through multiple small tricks in architecture and training that all together provide significant improvements in empirical performance [155].
@@
Multi-object tracking involves matching newly detected objects with the existing ones by their inter-frame positional and visual similarity information [18, 167]. ByteTrack improves the traditional Hungarian-algorithm-based matching paradigm to gather more comprehensive information [193]. BoT-SORT further incorporates advanced object re-identification modules and a refined Kalman filter for more accurate performances [5]. BoostTrack explores novel distance and shape similarity measurements to deal with ambiguity caused by unreliable detection results [142, 143]. Modern MOT pipelines are often implemented directly downstream to object detection in the tracking-by-detection paradigm.
@@
Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [8, 54]. Neural Social Physics model [185] incorporates learnable parameters into explicit physics models built on top of neural networks. SemanticFormer [147] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph. TrajNet++ [66], TDOR [53], and CASPNet++ [135] predict the distributions of future trajectories based on occupancy grid maps. Models like FRM [121] and PPT [86] decompose the prediction task by taking a multi-stage approach. Larger amounts of data comprised of multiple modalities and more comprehensive frameworks have shown increasing importance as is demonstrated by UniTraj++ [39]. Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
3.3 Real-time video analytics
@@
Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges. Video analytics requires the collection and processing of large volumes of video data, which can be resource-intensive and costly. Optimizing computation and network resource usage while maintaining or enhancing the accuracy of analytical results can be challenging. This challenge is further complicated by the need to adapt to varying network conditions, computational resources, and dynamic scene changes in real-time. Tab. 3 provides a comparative analysis of various approaches to address these challenges. The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions. This comparison provided in Tab. 3 highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
@@
Table 3: Comparison of various approaches and optimization objectives in video analytics.
@@
3.4 Use cases

### v2

# AI-Powered CPS-Enabled Urban Transportation Digital Twin: Methods and Applications

@@ Abstract
We present methods and applications for the development of digital twins (DT) for urban traffic management. While the majority of studies on the DT focus on its “eyes,” which is the emerging sensing and perception like object detection and tracking, what really distinguishes the DT from a traditional simulator lies in its “brain,” the prediction and decision making capabilities of extracting patterns and making informed decisions from what has been seen and perceived. In order to add value to urban transportation management, DTs need to be powered by artificial intelligence and complement with low-latency high-bandwidth sensing and networking technologies, in other words, cyberphysical systems (CPS). We will first review the DT pipeline enabled by CPS and propose our DT architecture deployed on a real-world testbed in New York City. This paper can be a pointer to help researchers and practitioners identify challenges and opportunities for the development of DTs; a bridge to initiate conversations across disciplines; and a road map to exploiting potentials of DTs for diverse urban transportation applications.
Index Terms:
Digital twin, AI, Urban traffic management
I Introduction
@@
Urban transportation systems are complex to model and simulate, due to heterogeneous road users (such as cars, pedestrians, cyclists, scooters) interacting in multimodal traffic environments consisting of public and private travel modes. With fast-changing traffic evolution in time and space, traffic simulation, if improperly calibrated, might produce traffic management strategies that largely deviate from the reality, potentially leading to suboptimal or even detrimental outcomes. With ubiquitous sensors in smart cities, it is the time to augment conventional traffic simulators, many of which were developed in an era when only “small data” became available. Emerging traffic sensors are expected to generate big volumes of data, transmitted via communication networks and processed on edge cloud computing with artificial intelligence (AI) for real-time traffic management. Such a transformation calls for the development of a new paradigm, namely, digital twin (DT), which will push the envelope in urban transportation management.
Literally, DT is the digital replica of a physical object or asset [1], where a digital world mirrors a physical world for real-time diagnosis, prognosis, and decision making. Recent years have seen a growing amount of studies on DT in various domains [1, 2, 3], including a sizable body of articles on vehicular DTs [4, 5, 6, 7, 8].
With recent explosive growth of literature on DTs, we would like to restrict the scope of this paper to applications in the urban setting, especially when vulnerable road users (VRU) (i.e., non-motorists such as pedestrian, bicyclists, other cyclists, or persons on personal conveyance [9]) are an integral part of the system and also potential users of the DT. We will primarily focus on use cases accounting for VRUs with improved traffic safety and efficiency.
The studies of DTs for urban traffic management, especially involving VRUs, are lacking, partly because the development of a DT for a system is non-trivial, particularly when involved with humans.
@@
This paper presents methods and applications for the development of DTs for urban transportation systems. We will depict a DT pipeline prototype, leveraging the architecture of cyberphysical systems and AI methods. We will propose a DT for real-time traffic monitoring and optimization, based on an existing physical testbed deployed in New York City (NYC) leveraging cutting-edge sensing, communication, computing, and AI-based automation. The overall contributions of this paper include: (1) introducing AI methods used in the DT pipeline; (2) exploring the architecture of transportation DTs and propose a prototype for reference; and (3) identifying gaps and directions.
@@
The rest of this paper is organized below. Section II introduces transportation DTs enabled by cyberphysical systems, and position this paper; Section III reviews the literature along the pipeline of a DT. Section IV demonstrates the architecture of our DT, building on a real-world testbed. Section V concludes our work, presents potential research directions and open questions.
II Preliminaries
@@
The definition of DT has evolved rapidly. Despite presenting its own version, articles share common elements and resemble certain characteristics [1, 10, 11]. In general, there is a physical world (aka. the physical) and a digital world (aka. the digital). The physical world evolves in time and space. To ensure that the physical system is run in a desired direction, it requires close monitoring, operation, and management. Thus, the role a digital world plays is to model and simulate the dynamics of the physical world in a synchronized fashion, so that the digital can also predict the future states of the physical precisely, which offers a ground for optimal decision making. The physical and the digital exchange data and information flows via a two-way communication. Specifically, the physical sends the data of its own state to the digital, and the digital feeds back the actuation signals to the physical. The actuation would trigger a change in the stage of the physical, and the updated state is sensed and sent back to the digital again. This iterative process runs between the physical and the digital as time unfolds. The sequential states of the physical should move towards a more desired state than without a DT. In other words, the ultimate goal of a DT development is to add values to the physical for improved safety and efficiency. Below, we first formalize transportation DT, and then discuss the relation between DT and cyberphyscial systems.
Definition II.1
@@
Transportation digital twin (T-DT) is a digital system integrating the pipeline from object detection and tracking, resource allocation, edge-cloud computing and communication, for online simulation, operation, control and management. It is updated online using continuously fed data collected from the physical and send control policies or issue warnings back to the physical, leveraging big data and AI tools. T-DTs are closed-loop with two-way communication, where data, information, and control signals are exchanged with the physical sequentially and reiteratively.
Definition II.2
@@
A transportation CPS interlinks physical and cyber layers, where the cyber layer consists of sensing, networking, computing, and traffic management application modules (see Fig. 1). The DT encloses the cyber layer, and relies on all the modules for two-way interaction with the physical. To enable the technological development of a DT, a physical testbed is needed as a platform for sensing, computing, experimentation, evaluation, as well as design constraints determination. In Sec. III, we will review the CPS technological enablers needed for the development of a DT, and examine the testbed used for our proposed DT in Sec. IV.
We would like to stress that, this paper aims to discuss how AI algorithms and the architecture of CPS contribute to the urban T-DT (UT-DT) pipeline. In contrast, there are highly cited survey papers on T-DT, which are more focused on general transportation scenarios and applications, while AI might not be the key focus. Tab. I outlines the comparison of a partial set of related work.
TABLE I: Comparison of survey papers on DT
Ref. Topic Focus
[4][5][6] Comprehensive review for vehicle mobility apps Vehicular technology driven pipeline
[13] Comprehensive review for traffic safety and mobility CPS pipeline highlighting communication
[14] Comprehensive review for operation and maintenance apps Broad travel modes
Ours Semi-technical review and position for urban traffic apps AI-Powered CPS pipeline
We first review various CPS and AI methods needed in the pipeline, and then present our T-DT instance. Accordingly, this paper is semi-survey, semi-technical. We employ such an organization, because we find that the publications on T-DT are normally segmented by different communities and journals, which could prevent researchers from understanding the entire pipeline, from upstream sensing and perception, to downstream transportation applications. For example, [15] on networking DT is primarily focused on the development of communication technology with evaluation on latency, even its use case is transportation. Transportation researchers, who hope to implement this system in real-world, have to seek more details about how traffic dynamics would impact the performance, which is unfortunately missing. This is likely because the authors belong to the society of communication and networking. Another example is that, [3] focused on vehicular DT, heavily rely on the foundational knowledge in CPS, which could be somewhat unfamiliar to many transportation researchers. Realizing such a gap in a large body of literature on T-DT, this paper aims to unify the knowledge by presenting a comprehensive summary upfront, and exemplify the pipeline following the summary. After all, the development of a T-DT calls for the interdisciplinary collaboration across transportation, electrical engineering, mechanical engineering, computer science, and human-machine interaction. To foster the readability of the paper, in the next section, we first offer a comprehensive state-of-the-art review on related work along the CPS pipeline.
III Related work on CPS-enabled DT
The development of CPS-enabled DT must engage with expertise in sensing, communication, computing, and human-centric perspectives, where AI methods are backbones.
Figure 1: Schematic diagram of a DT.
III-A Sensing and perception
@@
Sensors are the “eyes” (and “ears”) of a DT. Tab. II summarizes the pros and cons of each sensing technology, namely, mobile devices, on-board vehicles, and roadside infrastructure, for urban traffic applications. The former two are mobile sensors with wider spatial coverage but challenge in precision because of moving references, while the latter at fixed locations could face limited sensing ranges and coverage.
@@
TABLE II: Sensors for urban traffic applications (partially adapted from [16, 17])
@@
Camera Obstacle and lane detection 1. Maintains good resolution when the field of view is wide.
@@
2. Integrated with vehicles for behavior analysis [20].
@@
III-B Object detection and tracking
Object detection identifies and classifies objects within the environment using sensors like cameras, Lidar, and Radar. Object tracking is the process of monitoring the detected objects over time to determine their position and movement. Multi-object tracking is concerned with maintaining the identity of the objects and generating their trajectories. Trajectory prediction involves forecasting the future paths of detected and tracked objects. These tasks highly rely on training datasets for urban traffic scenes. Note that there is a much larger size of public datasets collected from on-board vehicles [25], but fewer from other sensor types. Here we thus summarize commonly used and emerging datasets from non-vehicle sensors in Tab. III.
TABLE III: Public datasets for urban traffic scenes (On-board vehicle datasets can be found in [25].)
@@ III-B Object detection and tracking
WILDTRACK [30] Multi-Object Tracking RGB cameras University campus in Zurich
Aerial VisDrone [31], NGSim [32], highD [33], rounD [34], DOTA [35], CitySim [36] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Urban spaces in China and Aachen, intersections in California and Florida
MOT Challenge [37], MOTS20 [38], UAVDT [39] Aerial object detection and tracking, trajectory forecasting Drone-based RGB cameras Various Indoor and Outdoor Scenes
@@
Synthetic CARLA [42, 43, 44] Autonomous-driving object detection and segmentation RGB cameras, LiDAR Urban European/North American environments
GTAV [45], Synscapes [46], UrbanSyn [47] Object detection and segmentation RGB cameras, LiDAR Urban European/North American environments
MOTSynth [48] Multi-object tracking RGB cameras Urban European/North American environments
MatrixCity [49] Neural-rendering benchmarking (vehicle/pedestrian-free) RGB cameras Synthetic city environment
Boundless [50] Object detection and segmentation with UE5-synthesised data RGB cameras Synthetic urban environments
Object detection has been studied extensively for urban applications. A large number of studies focus on low-altitude vehicle and pedestrian detection [51, 52]. Many focus on high-altitude aerial environments, where small object detection becomes an important challenge [31, 35, 33, 36]. Single-stage object detectors, following the original single shot multibox detector [53] and You Look Only Once (YOLO) [54] architectures, have become popular due to their real-time deployment capabilities. In the last few years, transformer-based object detection approaches, competitive with YOLO models, have emerged as the state-of-the-art in object detection when designed to be deployed in the real-time setting [55, 56, 57, 58]. Recent progress in YOLO object detection performance has been enabled through multiple small tricks in architecture and training that all together provide significant improvements in empirical performance [59].
@@
Multi-object tracking involves matching newly detected objects with the existing ones by their inter-frame positional and visual similarity information [72, 73]. ByteTrack improves the traditional Hungarian-algorithm-based matching paradigm to gather more comprehensive information [74]. BoT-SORT further incorporates advanced object re-identification modules and a refined Kalman filter for more accurate performances [75]. BoostTrack explores novel distance and shape similarity measurements to deal with ambiguity caused by unreliable detection results [76, 77].
@@
Predicting future trajectories of detected objects is often a crucial part of safety-critical applications. Numerous deep neural network models have emerged as competitive candidates for trajectory prediction over the past few years. Majority of modern architectures for predicting future trajectories of detected objects adopt Recurrent Neural Networks which is responsible for predicting future object positions based on their historical coordinates, together with generative components which handles the variation and flexibility in social interactions [78, 79]. Neural Social Physics model [80] incorporates learnable parameters into explicit physics models built on top of neural networks. SemanticFormer [81] seeks more structural and humanized environment understanding by constructing a semantic knowledge graph. TrajNet++ [82], TDOR [83], and CASPNet++ [84] predict the distributions of future trajectories based on occupancy grid maps. Models like FRM [85] and PPT [86] decompose the prediction task by taking a multi-stage approach. Larger amounts of data comprised of multiple modalities and more comprehensive frameworks have shown increasing importance as is demonstrated by UniTraj++ [87]. Unlike the tracking algorithms, specific training or fine-tuning is often required before the deployment of trajectory forecasting models to unseen scenarios.
III-C Real-time video analytics
@@
Developing end-to-end real-time video analytics systems on a large scale for time-sensitive and safety-critical traffic and crowd management applications presents challenges. Video analytics requires the collection and processing of large volumes of video data, which can be resource-intensive and costly. Optimizing computation and network resource usage while maintaining or enhancing the accuracy of analytical results can be challenging. This challenge is further complicated by the need to adapt to varying network conditions, computational resources, and dynamic scene changes in real-time. Tab. IV provides a comparative analysis of various approaches to address these challenges. The approaches differ in their focus–some prioritize reducing latency and resource consumption, while others emphasize maintaining or enhancing accuracy, especially under constrained conditions. This comparison provided in Tab. IV highlights the trade-offs inherent in real-time video analytics and emphasizes different optimization methods to balance throughput, accuracy, energy consumption, and computational efficiency across diverse deployment scenarios, including edge devices, cloud platforms, and hybrid environments.
@@
TABLE IV: Comparison of various approaches and optimization objectives in video analytics.
@@
III-D Communication and networking
III-D1 Real-time requirements and low latency targets
Smart city applications can be grouped according to their latency requirements. Many, if not all, pedestrian-associated application (facilitated by pedestrian detection/observations and message notifications) are likely to expect the round trip delay in the range of a couple of seconds. Such latency can be supported by contemporary cameras, communication protocols and inferencing engines. Applications which would attempt to close the observation/notification loop for vehicles moving at about 10 km/h may expect latency in tens of millisecond. Using conventional video compression, RTP/RTSP streaming and edge computing is inadequate to support such latency. This presents the opportunity to pursue novel engineering solutions and research problems.
III-D2 Communication techniques and protocols
Tab. V summarizes the key characteristics of commonly used IoT communication protocols, Message Queuing Telemetry Transport (MQTT) [120], Constrained Application Protocol (CoAP) [121], and Hypertext Transfer Protocol (HTTP) [122]. Among them, MQTT emerges as a practical choice, due to its combination of low latency, high scalability, and reliable delivery mechanisms. Its lightweight publish/subscribe model, Quality of Service (QoS) guarantees, and session management make it well suited for real-time data exchange in dynamic, safety-critical environments like urban transportation systems [123, 124].
TABLE V: Communication protocol comparison for real-time digital twin systems.
Protocol Latency Scalability Reliability Best Suited For
MQTT Persistent TCP connection delivers low latency for ongoing messaging Highly scalable with a broker-based publish/subscribe model supporting many-to-many communication across thousands of devices Guarantees reliable delivery with configurable acknowledgment levels, exactly-once delivery via a four-step handshake, persistent sessions Large-scale IoT, real-time sensor/actuator networks
CoAP UDP-based, very low latency when network is stable, but performance degrades with packet loss Limited scalability due to client-server request/response model. Multicast is possible but complex and unreliable at scale Basic two-way acknowledgment, no exactly-once delivery guarantee, no session continuity Resource-limited devices that send small, infrequent data
HTTP Higher latency due to new TCP/TLS handshake and headers per request Limited scalability due to client-server request/response model and verbose ASCII headers Relies on TCP for delivery, with no application-level acknowledgment, retries, or session handling Web APIs, periodic data transfer, backend integration
IV Proposed DT pipeline
Figure 2: Architecture of the proposed DT pipeline: Urban Transportation DT (UT-DT).
In this section, we present a DT architecture for UT-DT, based on the sensing/communication/computing testbed deployed in NYC (Fig. 2). The proposed architecture is enabled by cameras and LiDARs, high speed communications, and edge cloud computing.
Figure 3: Diagram of intersection safety warning use case.
IV-A Physical Infrastructure
Pilot experiments are executed at the signalized intersection of Amsterdam Avenue (major) and 120th Street (minor) near Columbia campus in NYC. The road geometry and traffic statistics are summarized in Fig. 4.
Figure 4: Traffic statistics for NYC’s intersection of 120th St. & Amsterdam Ave.
IV-A1 Technological enabler
The physical functionality of the proposed UT-DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment”) [125], developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125]. It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications. Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
At the intersection of Amsterdam Avenue and 120th Street, cameras, LiDAR, and wireless sensing and communication nodes are deployed. The cameras employ H264 encoding with an I-frame interval set to 10 frames. The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU. The LiDAR is a Velodyne VLP-32C with up to 200m range.
IV-B Digital twin
[…]
(a) COSMOS in CARLA
(b) COSMOS in SUMO
(c) Boundless with snow
(d) Boundless with rain
Figure 5: (Top) Carla-SUMO simulator; (Bottom) Qualitative examples from Boundless. Bounding boxes for vehicles are shown in blue, and pedestrians are shown in green.
IV-C Use cases